OpenStudy (anonymous):

Find the maximum value of the product of two numbers x and x-10 which add up to 10.

OpenStudy (anonymous):

5 by pure reason

OpenStudy (anonymous):

so how did you calcutlate it?

OpenStudy (anonymous):

i assume you mean x and 10 - x since they add to ten. each number is five and the product is 25

OpenStudy (anonymous):

oh yea, thats right! thanks for your help mate :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if you have to write something for your teacher, write the product is \[x(10-x)=10x-x^2\] which will have a maximum value at the vertex which is \[-\frac{b}{2a}=\frac{10}{2}=5\]

OpenStudy (anonymous):

okay, but how did you get the second part?

OpenStudy (mertsj):

5 is the x coordinate of the vertex. To get the y coordinate plug x into 10x - x^2

OpenStudy (anonymous):

okay, so a=1 and b=10? do you just ignore the negative?

OpenStudy (anonymous):

no because \[a=-1\] and there is a -1 in the 'formula' which is why it is postive

OpenStudy (anonymous):

anyway to heck with the formula. you have two numbers and this setup is completely symmetric, but which i mean you cannot tell the numbers apart. x and 10 - x or 10 - x and x so clearly it is the biggest when they are the same

OpenStudy (anonymous):

oh i see! that makes sense now, thankss! sorry to irritate you lol

OpenStudy (anonymous):

no irritation, no problem. i only wrote that vertex thing in case you had to hand something in to your teacher